Solve the following system by substitution:
Solve the following system by substitution:

y =x +9
x +9y =21
If the equations of the system are dependent, or if a system is inconsistent, so indicate.
In those cases enter dependent or inconsistent. Otherwise, enter your answer as an ordered pair.


Answer by Nate(3500)

y =x +9
x +9y =21
x +9(x+9) =21
x + 9x + 81 = 21
10x=-60
x=-6
--------------------
y =x +9
y=(-6)+9
y=3
--------------------
these systems have different slopes, so they are perpendicular (independent;consistent)
the point they intersect is: (-6,3)